40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 18 18:33:57 2014, ericq, Update, Green Locking, Yarm Green PDH yLoop.pdfcrap.pdf
    Reply  Tue Aug 19 01:01:36 2014, Jenne, Rana, Update, Green Locking, Yarm Green PDH YPDH_OLG.pdfYPDH_noise.pdf
       Reply  Tue Aug 19 18:32:40 2014, ericq, Update, Green Locking, Yarm Green PDH 
       Reply  Tue Aug 19 23:11:15 2014, Jenne, Update, Green Locking, Yarm Green PDH ErrAndCtrlSpectra_VoltsPerRtHz.png
       Reply  Wed Aug 20 02:38:41 2014, Jenne, Update, Green Locking, Yarm Green PDH - requirement NoiseConsideration.pdf
          Reply  Wed Aug 20 18:05:18 2014, Jenne, Update, Green Locking, Yarm Green PDH - requirement YPDH_noise.pdf
       Reply  Wed Aug 20 04:09:21 2014, ericq, Update, Green Locking, Xarm Green PDH Xbodes.pdfXspectra.pdfscopeSweep.jpg
          Reply  Wed Aug 20 16:10:43 2014, ericq, Update, Green Locking, Xarm Green PDH XspectraCombined.pdf
             Reply  Wed Aug 20 21:09:16 2014, ericq, Update, Green Locking, Xarm Green PDH scopeSweep.jpgXspectra.pdfYspectra.pdfXYcomp.pdf
Message ID: 10408     Entry time: Tue Aug 19 01:01:36 2014     In reply to: 10407     Reply to this: 10409   10411   10412   10413
Author: Jenne, Rana 
Type: Update 
Category: Green Locking 
Subject: Yarm Green PDH 

[ Rana, Jenne]

We remeasured the Yend PDH box.

When we first started, the green couldn't hold lock to the arm - it kept flickering between modes.  Changing the gain of the PDH box (from 7.5 to 6.0) helped.

We measured a calibration, from our injection point to our measurement point.

The concept was that we'd take the mixer output, and put that into an SR560, and put the swept sine injection into the other input port of the '560, and use A-B.  So, for this calibration, we left A unplugged, and just had the RF out of the 4395 going to input B of the '560.  The 600 Ohm output of the '560 went to the error point input on the PDH box (during normal operation the mixer output is connected directly to the error point input).  The SR560 was set to gain of 1, no filtering.  I don't recall if we were using high range or low noise, but we tried both and didn't really see a difference between them.

We had the 4395 take that calibration out, and then we measured the closed loop gain up to 1 MHz. (Same measurement setup as above, but we connected the mixer out to the input of the SR560 to close the loop, and made sure we were locked on a TEM00 green mode.) Rana used an ipython notebook to infer the open loop gain from our measurement.  Our conclusion is that we don't have nearly enough gain margin in our loop.  We found the PDH box gain knob at 7.5, and we turned it down to 6.0, but the loop is still pretty borderline. We used the high impedance active probe to measure the error point monitor, since we aren't sure that that point can drive a 50 Ohm load.

YPDH_OLG.pdf

We also measured the error point spectra and the control point spectra.  Unfortunately, the saved data from the analyzer (no matter what is on the screen) comes out in spectrum, not spectral density.  So, we need to check our conversion, but right now to get from Watts power to Volts, we do sqrt(50 ohm * data).  We then need to get to spectral density, and right now we're just dividing by the square root of the bandwith that is reported in the .par file. This last step is the one we want to especially check, by perhaps putting some known amount of noise (from an SR785?) into the 4395, and checking that our calibration math returns the expected noise spectrum.

What still needs to be done is to calibrate this into Hz/rtHz.  To do this, we were thinking that we should look at the error point on a 'scope while the cavity is flashing.

Anyhow, here is the uncalibrated error point spectrum.  Purple is a measurement up to 30kHz, with 30Hz bandwidth.  Blue is a measurement up to 300kHz with 300Hz bandwidth.  The gain peaking schmutz above 10kHz sucks, and we'd like to get rid of it.  We also see the same peak at ~150kHz that Q saw earlier today.  We were using the high impedance probe here too.

YPDH_noise.pdf

 We have the data for the control point (all the data files are in /users/jenne/ALS/PDHloops/Yend_18Aug2014), but we haven't plotted it yet.

Things that need doing:

* (JCD) Think about this box's purpose in life.  What kind of gain do we need?  Do we need more / less than we're currently getting? NPRO freq noise is 1/f and is 10kHz/rtHz at 1Hz (this is from a plot of an iLIGO NPRO from Rana's thesis, but it's probably similar). Talk to Kiwamu; the noise budget in the paper seems to indicate that we had some kind of boost on or something.  Also, if we need much more gain than we already have, we'll definitely need a different box, maybe the PDH2 box that they have over in WBridge.

* (EQ, priority 1) Measure and calibrate error point noise down to lower freq for both arms.  What could we win by putting in a boost? If the residual noise is high, maybe the laser isn't good at following arm, so beatnote isn't good length info for the arm, and we can't succeed.

* (EQ, priority 2) Measure TF of PDH box, and a separate measurement of the Pomona box that is between the mixer and the error point - is that eating a bunch of phase?  It's already an LC circuit which is good, but do we really want a 120kHz lowpass when our modulation frequency is roughly 200kHz?  Ask ChrisW - he worked on one of these with Dmass.

* (EQ, priority 2ish) Measure TF of Xend PDH loop (unless you already have one, up to ~1MHz).

* (JCD) Make DCC tree leaf for PDH box #17.  Take photos of box.

ELOG V3.1.3-